OpenStudy (anonymous):

2 0 doctor_will617 A linear function L is defined so that L(5)=7 and L(-1) =25. Find a formula for L of the form L(x)=mx+b. Then graph L, showing both intercepts and labeling them with their coordinates

OpenStudy (anonymous):

Ok well lets figure out what the points are first: when x=5, L(5)=7 so (5,7) when x= -1, L(-1) = 25 so (-1,25) now we just have to find the slope (m). Slope is just the Rise over Run or (L(-1) - L(5)) / (-1 - 5) which now gives us: m = (25-7) / (-1-5) = 8/(-6) = -4/3 so you have L(x) = -4/3 (x) + b plug in one of your points, lets do (5,7) 7 = (-4/3)(5) + b 7= (-20/3) + b isolate the (b) and you get b = 1/3 so you're equation for the line is: L(x) = (-4/3)(x) + (1/3)

OpenStudy (anonymous):

isn't it 18/-6=-3

OpenStudy (anonymous):

for the slope

OpenStudy (anonymous):

HAHA OH MAN you're right it should be -3

OpenStudy (anonymous):

but you get the gist right?

OpenStudy (anonymous):

yup thanx